The second angle of a triangle measures 3 times the first and the third angle measures 52 degrees more than the sum of the measures of the other two.What are the angles?

Answer by stanbon(75887)

The second angle of a triangle measures 3 times the first and the third angle measures 52 degrees more than the sum of the measures of the other two.What are the angles?
--------------
1st: x
2nd: 3x
3rd: x + 3x + 52
-----------
Equation:
x + 3x + x + 3x + 52 = 180
8x + 52 = 180
8x = 128
1st: x = 16 degrees
2nd: 3x = 48 degrees
3rd: 4x+52 = 116 degrees
-----------------------------
Cheers,
Stan H.
--------